OpenStudy (anonymous):

Find the extreme point(s) of the function f(x) = 0.25^4 + 3x^3 -18x^2 +10 and classify them

OpenStudy (amistre64):

local extrema or global?

OpenStudy (anonymous):

Not quite sure

OpenStudy (amistre64):

its a quarctic...probably spelled it wrong... that will look like some form of a "W"

OpenStudy (amistre64):

you know how to find derivatives?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

thats all this is, get the 1st derivative and set it to zero

OpenStudy (anonymous):

Thanks much!

OpenStudy (amistre64):

f'(x) = x^3 + 9x^2 -36x

OpenStudy (amistre64):

f'(x) = x(x^2 +9x -36) = 0 at x=0 is an obvious one..and it should be a MAX

OpenStudy (amistre64):

3,12 (x+12)(x-3)x = -12 and x = 3 are 2 more

OpenStudy (anonymous):

how do you know its a max?, x=0 is a max